# SVInvNet: A Densely Connected Encoder-Decoder Architecture for Seismic Velocity Inversion

### v1

@@ V Conclusion
In this study, we introduce a novel end-to-end deep model, SVInvNet, for seismic velocity inversion. In order to enhance the capacity of SVInvNet in learning and estimating more complex velocity models, we abstained from merely increasing the channel number of CNN layers (i.e. number of feature planes), which cause an excessive increase in the number of parameters. Instead, we devised a network architecture characterized by the multiple connections among its layers with only 4M parameters. This design facilitated the efficient flow of information between the input and the corresponding ground truth. Based on the outcomes derived from both noisy and noiseless datasets, SVInvNet demonstrates proficiency in accurately estimating velocity models characterized by multilayer structures, fault, and salt dome. Furthermore, it is observed that increasing the quantity of training dataset correlates with a reduction in loss function in the testing phase. The utilization of a test dataset size at least twice that of the training datasets reveals the comprehensiveness of the results produced by the trained SVInvNet.

### v2

@@ V Conclusion
In this study, we introduce a new end-to-end deep model, SVInvNet, for seismic velocity inversion. In order to enhance the capacity of SVInvNet in learning and estimating more complex velocity models, we abstained from merely increasing the channel number of CNN layers (i.e. number of feature planes), which cause an excessive increase in the number of parameters. Instead, we devised a network architecture characterized by the multiple connections among its layers with only 4M parameters. This design facilitated the efficient flow of information between the input and the corresponding ground truth. Based on the outcomes derived from both noisy and noiseless datasets, SVInvNet demonstrates proficiency in accurately estimating velocity models characterized by multilayer structures, fault, and salt dome. Furthermore, it is observed that increasing the quantity of training dataset correlates with a reduction in loss function in the testing phase. The utilization of a test dataset size at least twice that of the training datasets reveals the comprehensiveness of the results produced by the trained SVInvNet.